### v3

@@ 3 Application to complex SDP relaxations for CPOPs
(7) $[A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}},&\text{if }({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
@@
Therefore, one can rewrite ($-{\mathbb{C}}$) as follows:
($-{\mathbb{C}}$’) $\begin{cases}\inf\limits_{y}&\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&({\mathscr{A}}^{i})^{*}(y)\succeq 0,\quad i=0,1,\ldots,t,\\ &y_{\mathbf{0},\mathbf{0}}=1,\end{cases}$
@@
($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Note that we have used the Kronecker delta $\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}$ in ($-{\mathbb{C}}$).
Let us from now on fix any order “$<$” on ${\mathbb{N}}^{s}$.
@@
($-{\mathbb{C}}$) is equivalent to the following complex SDP:
($-{\mathbb{C}}$’) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
@@
Let ${\mathscr{A}}^{i}_{R},{\mathscr{A}}^{i}_{I}$ be the real linear operator associated with ${\mathscr{A}}^{i}$ which are defined in a similar way as (4) and (5). With $H^{i}=H^{i}_{R}+H^{i}_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$, ($-{\mathbb{C}}$) is equivalent to the following real SDP by using the equivalent condition (1):
($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{\lambda,Y^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &Y^{i}=\begin{bmatrix}H^{i}_{R}&-H^{i}_{I}\\ H^{i}_{I}&H^{i}_{R}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
On the other hand, by invoking Theorem 2.3, we obtain another equivalent real SDP conversion of ($-{\mathbb{C}}$):
@@
($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad{\boldsymbol{\beta}}\neq{\boldsymbol{\gamma}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
@@
[…]
where we have used fact that $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$ and $[(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}$. It follows that $[{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\rangle=0$.
@@
[…]
where we have used fact that ${\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}})={\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+\overline{g}^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})=0$ and $X^{i}_{1},X^{i}_{2}$ are symmetric. It follows that $[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\rangle=0$.
@@
($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
@@ 4 Numerical experiments
In this section, we benchmark the performance of the two real reformulations for complex SDPs using the software TSSOS 1.2.1333TSSOS is freely available at https://github.com/wangjie212/TSSOS. in which MOSEK 10.0 [1] is employed as an SDP solver with default settings. For comparison, we also include the results of directly solving complex SDPs obtained with Hypatia 0.8.1 [6] in Sections 4.14.3. All numerical experiments were performed on a desktop computer with Intel(R) Core(TM) i9-10900 CPU@2.80GHz and 64G RAM. When presenting the results, the column labelled by ‘opt’ records the optimum and the column labelled by ‘time’ records running time in seconds. Moreover, the symbol ‘-’ means the SDP solver runs out of memory, and the symbol ‘*’ means running time exceeds 10000 seconds.
@@ 4.1 Minimizing a random complex quartic polynomial over the unit sphere
Our first example is to minimize a complex quartic polynomial over the unit sphere:
@@
where $[{\mathbf{z}}]_{2}$ is the column vector of monomials in ${\mathbf{z}}$ up to degree two and $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the standard normal distribution.
We approach (10) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 2. From the table, we see that the reformulation ($-{\mathbb{R}}$) is several ($2\sim 7$) times as fast as the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows. On the other hand, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
@@
$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)

### v4

@@ 3 Application to complex SDP relaxations for CPOPs
(7) $A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}},&\text{if }({\boldsymbol{\beta}}^{\prime}+{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime}+{\boldsymbol{\gamma}}^{\prime\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$
@@
Therefore, one can rewrite (Mom-$\mathbb{C}$}) as follows:
(Mom-$\mathbb{C}$}’) $\begin{cases}\inf\limits_{y}&\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&({\mathscr{A}}^{i})^{*}(y)\succeq 0,\quad i=0,1,\ldots,t,\\ &y_{\mathbf{0},\mathbf{0}}=1,\end{cases}$
@@
(HSOS-$\mathbb{C}$}) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
Note that we have used the Kronecker delta $\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}$ in (HSOS-$\mathbb{C}$}).
Let us from now on fix any order “<<” on ${\mathbb{N}}^{s}$.
@@
(HSOS-$\mathbb{C}$}) is equivalent to the following complex SDP:
(HSOS-$\mathbb{C}$}’) $\begin{cases}\sup\limits_{\lambda,H^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}[{\mathscr{A}}^{i}(H^{i})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &H^{i}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
@@
Let ${\mathscr{A}}^{i}_{R}$ and ${\mathscr{A}}^{i}_{I}$ be the real linear operators associated with ${\mathscr{A}}^{i}$ which are defined in a similar way as (4) and (5). With $H^{i}=H^{i}_{R}+H^{i}_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$, (HSOS-$\mathbb{C}$}) is equivalent to the following real SDP by using the equivalent condition (1):
(HSOS-$\mathbb{R}$}) $\begin{cases}\sup\limits_{\lambda,Y^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(H^{i}_{I})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(H^{i}_{R})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &Y^{i}=\begin{bmatrix}H^{i}_{R}&-H^{i}_{I}\\ H^{i}_{I}&H^{i}_{R}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
On the other hand, by invoking Theorem 2.3, we obtain another equivalent real SDP conversion of (HSOS-$\mathbb{C}$}):
@@
(HSOS-$\mathbb{R}$}’) $\begin{cases}\sup\limits_{\lambda,X^{i}}&\lambda\\ \,\rm{s.t.}&\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}+[{\mathscr{A}}^{i}_{I}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)+\delta_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\,(\mathbf{0},\mathbf{0})}\lambda=[b_{R}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\\ &\sum_{i=0}^{t}\left([{\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}-[{\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\right)=[b_{I}]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad{\boldsymbol{\beta}}\neq{\boldsymbol{\gamma}},\\ &{\boldsymbol{\beta}}\leq{\boldsymbol{\gamma}},\quad({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{s}_{d}\times{\mathbb{N}}^{s}_{d},\\ &X^{i}=\begin{bmatrix}X^{i}_{1}&(X^{i}_{3})^{\intercal}\\ X^{i}_{3}&X^{i}_{2}\end{bmatrix}\succeq 0,\quad i=0,1,\ldots,t.\end{cases}$
@@
$=$  $\,0,$
where we have used fact that $(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}$ and $(X^{i}_{3})^{\intercal}]_{{\boldsymbol{\gamma}}^{\prime},{\boldsymbol{\beta}}^{\prime}}=[X^{i}_{3}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}$. It follows that ${\mathscr{A}}^{i}_{R}(X^{i}_{3}-(X^{i}_{3})^{\intercal})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{R},X^{i}_{3}-(X^{i}_{3})^{\intercal}\rangle=0$.
@@
$=$  $\,0,$
where we have used fact that ${\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+g^{i}_{{\boldsymbol{\gamma}}^{\prime\prime},{\boldsymbol{\beta}}^{\prime\prime}})={\mathcal{I}}(g^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}}+\overline{g}^{i}_{{\boldsymbol{\beta}}^{\prime\prime},{\boldsymbol{\gamma}}^{\prime\prime}})=0$ and that $X^{i}_{1},X^{i}_{2}$ are symmetric. It follows that ${\mathscr{A}}^{i}_{I}(X^{i}_{1}+X^{i}_{2})]_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}}=\langle(A^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\beta}}})_{I},X^{i}_{1}+X^{i}_{2}\rangle=0$.
@@
(HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$})
@@ 4 Numerical experiments
In this section, we benchmark the performance of the two real reformulations for complex SDPs using the software TSSOS 1.2.1333TSSOS is freely available at https://github.com/wangjie212/TSSOS. in which MOSEK 10.0 [1] is employed as an SDP solver with default settings. For comparison, we also include the results of directly solving complex SDPs obtained with Hypatia 0.8.1 [8] in Sections 4.14.4. We emphasis that the purpose of numerical experiments is to compare different formulations of complex SDPs and a comparison between solvers is not our focus in this paper. All numerical experiments were performed on a desktop computer with an Intel(R) Core(TM) i9-10900 CPU@2.80GHz and 64G RAM. When presenting the results, the column labelled by ‘opt’ records the optimum and the column labelled by ‘time’ records computational time in seconds. Moreover, the symbol ‘-’ means the SDP solver runs out of memory, and the symbol ‘*’ means computational time exceeds 10000 seconds.
4.1 The complex matrix completion problem
The complex matrix completion problem seeks to recover a low-rank matrix $M\in{\mathbb{C}}^{s\times t}$ from a subset of entries $\{M_{ij}\}_{(i,j)\in\Omega}$ [6, 7], which could be tackled via the convex optimization problem:
(10) $\begin{cases}\inf\limits_{Z\in{\mathbb{C}}^{s\times t}}&\|Z\|_{*}\\ \,\,\,\,\,\,\rm{s.t.}&Z_{ij}=M_{ij},\quad\forall(i,j)\in\Omega,\end{cases}$
where $\|Z\|_{*}\coloneqq\hbox{\rm{Tr}}(Z^{{\mathrm{H}}}Z)^{\frac{1}{2}}$ is the nuclear norm of $Z$. One can equivalently cast (10) as a complex SDP:
(11) $\begin{cases}\inf\limits_{X\in{\mathrm{H}}_{n}}&\hbox{\rm{Tr}}(X)\\ \,\,\,\rm{s.t.}&Z_{ij}=M_{ij},\quad\forall(i,j)\in\Omega,\\ &X=\begin{bmatrix}U&Z^{{\mathrm{H}}}\\ Z&V\end{bmatrix}\succeq 0.\end{cases}$
In this subsection, we generate random instances of the complex matrix completion problem (10) as follows: 1) select $\Omega\subseteq[s]\times[s]$ uniformly at random from all subsets with cardinality $10s$ for $s\in\{20,30,40,50,60,70\}$; 2) set $M=M_{1}M_{2}^{{\mathrm{H}}}$ with the real and imaginary parts of entries of $M_{1}\in{\mathbb{C}}^{s\times 5}$ and $M_{2}\in{\mathbb{C}}^{s\times 5}$ being selected i.i.d. from the standard normal distribution. The related results are shown in Table 2. From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is $1\sim 3$ magnitudes faster than the reformulation (HSOS-$\mathbb{R}$}).
Table 2: The complex matrix completion problem.
$s$ (PSDP-$\mathbb{R}$}) (PSDP-$\mathbb{R}$}) (PSDP-$\mathbb{C}$})
$m_{\rm{sdp}}$ opt time $m_{\rm{sdp}}$ opt time opt time
20 1948 315.882 0.62 308 315.882 0.06 315.882 3.46
30 4168 456.591 5.38 508 456.591 0.11 456.591 31.1
40 7160 670.611 62.7 680 670.611 0.37 670.611 153
50 11010 861.510 175 910 861.510 0.77 861.510 626
60 15602 925.590 573 1082 925.590 1.07 925.590 1899
70 21056 1108.93 1337 1316 1108.93 1.81 1108.93 4870
@@ 4.2 Minimizing a random complex quartic polynomial over the unit sphere
Our second example is to minimize a complex quartic polynomial over the unit sphere:
@@
where ${\mathbf{z}}]_{2}$ is the column vector of monomials in $\mathbf{z}$} up to degree two and $Q\in{\mathrm{H}}^{|[{\mathbf{z}}]_{2}|}$ is a random Hermitian matrix whose entries are selected with respect to the standard normal distribution.
We approach (12) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 3. From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is several ($2\sim 7$) times as fast as the reformulation (HSOS-$\mathbb{R}$}), and the speedup becomes more significant as the SDP size grows.
@@
$s$ $d$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})